# A mathematical theory of topological invariants of quantum spin systems

### v2

@@ 6.1 The commutator class
Even if ${\mathfrak{M}}$ is acyclic, the DGLA ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$ is not necessarily acyclic. We would like to construct an obstruction to finding an element ${\mathsf{p}}\in{\mathfrak{M}}_{-1}$ which satisfies $\partial{\mathsf{p}}={\mathsf{B}}$ and $[{\mathsf{p}},{\mathsf{p}}]=0$.
Definition 6.2.
@@
$\partial{\mathsf{p}}+\frac{1}{2}[{\mathsf{p}},{\mathsf{p}}]={\mathsf{B}}.$
@@
Let $({\mathfrak{M}},{\mathsf{B}})$ be pronilpotent pointed DGLA and ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$. Then $[{\mathsf{p}},{\mathsf{p}}]$ is a cycle of the DGLA ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$.
Proposition 6.1.
Let $({\mathfrak{M}},{\mathsf{B}})$ be an acyclic pointed DGLA which is a limit of an inverse system of nilpotent acyclic pointed DGLAs $({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $N\in{\mathbb{N}}$. Assume further that the structure morphisms $r_{N,N-1}:{\mathfrak{M}}_{N}\rightarrow{\mathfrak{M}}_{N-1}$ are surjective and ${\mathfrak{j}}_{N}=\ker r_{N,N-1}$ is central in ${\mathfrak{M}}_{N}$. Finally, assume $[{\mathfrak{M}}_{1},{\mathfrak{M}}_{1}]=0$. Then ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is non-empty. Furthermore, the homology class of $[{\mathsf{p}},{\mathsf{p}}]$ in ${\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}$ for ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is independent of the choice of ${\mathsf{p}}$.
A proof of this result can be found in Appendix B. It uses some results from deformation theory. We will call the homology class of $[{\mathsf{p}},{\mathsf{p}}]$ the commutator class of the acyclic pointed DGLA $({\mathfrak{M}},{\mathsf{B}})$, for lack of a better name, and denote it ${\mathbf{com}}({\mathfrak{M}},{\mathsf{B}})$. The assignment $({\mathfrak{M}},{\mathsf{B}})\mapsto(H_{\bullet}({\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}),{\mathbf{com}}({\mathfrak{M}},{\mathsf{B}}))$ is a functor from the full sub-category of the category of acyclic pointed DGLAs satisfying the conditions of Proposition 6.1, to the category ${{\mathsf{pVect}}_{\mathbb{Z}}}$ of pointed graded vector spaces.
The conditions of Prop. 6.1 apply to any pointed DGLA which is the value of the Čech functor on a graded local Lie algebra ${\mathfrak{F}}^{\bf G}$ over some $(X,\leq)$, where ${\mathfrak{F}}$ is a local Lie algebra equipped with a smooth infinitesimally inner $G$-action. Indeed, along with the graded algebra $A=\prod_{k=1}^{\infty}{\rm Sym}^{k}{\mathfrak{g}}^{*}[-2]$ used in the construction of ${\mathfrak{F}}^{\bf G}$ one can consider its quotient by the ideal $J_{N}=\prod_{k=N+1}^{\infty}{\rm Sym}^{k}{\mathfrak{g}}^{*}[-2]$. Replacing $A$ with the nilpotent graded algebra $A/J_{N}$ throughout, for any cover ${\mathfrak{U}}$ of the terminal object $T$ one gets a sequence of nilpotent acyclic pointed DGLAs labeled by $N\in{\mathbb{N}}$. They assemble into an inverse system in an obvious manner, and its limit is the acyclic pointed DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},T;{\mathfrak{F}}^{\bf G}).$ It is easy to see that the remaining conditions of Prop. 6.1 are also satisfied. In particular, Prop. 6.1 applies to the pointed DGLAs associated the graded local Lie algebras ${\mathfrak{D}}^{\psi,\bf G}_{al}$ and ${{\mathfrak{g}}^{\bf G}_{al}}$ and any CS cover of ${\mathbb{R}}^{n}$.
Proposition 6.2.
@@
Let ${\mathfrak{V}}=\{V_{j}\}_{j\in J}$ be a refinement of ${\mathfrak{U}}$. By definition, there exists a map $\phi:J\rightarrow I$ with $V_{j}\subset U_{\phi(j)}.$ According to Prop. 2.2, there is a map from ${\mathfrak{M}}({\mathfrak{V}})$ to ${\mathfrak{M}}({\mathfrak{U}})$. As ${\mathsf{B}}$ is unaffected by $\phi_{*}$, we deduce from ${\mathsf{p}}^{\mathfrak{V}}\in{\rm MC}({\mathfrak{M}}({\mathfrak{V}}),{\mathsf{B}})$ that $\phi_{*}{\mathsf{p}}^{\mathfrak{V}}\in{\rm MC}({\mathfrak{M}}({\mathfrak{U}}),{\mathsf{B}})$. ∎
Example 6.1.
Let $G$ be a compact Lie group, ${\mathfrak{g}}$ be its Lie algebra, and ${{\mathfrak{g}}_{al}}$ be the pointed local Lie algebra over ${\mathcal{C}\mathcal{S}}_{n}$ of Examples (5.2) and (5.5). The distinguished central cycle in $C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n},{{\mathfrak{g}}^{\bf G}_{al}})$ is the constant function on $\Lambda$ with value ${\rm id}_{\mathfrak{g}}$ Here ${\rm id}_{\mathfrak{g}}$ is regarded as a $G$-invariant element of ${{\mathfrak{g}}_{al}}({\mathbb{R}}^{n})\otimes{\mathfrak{g}}^{*}[-2]$. For any cover ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ one can construct a twisted MC element ${\mathsf{q}}$ as follows. Pick $r>0$ large enough so that the interiors of $U_{i}^{r}$, $i\in I$, cover ${\mathbb{R}}^{n}$ in the usual sense and pick a partition of unity $\chi_{i}$, $i\in I$, subordinate to this open cover. For any $j\in\Lambda$ and any $i\in I$ let ${\mathsf{q}}_{i}(j)=\chi_{i}(j){\rm id}_{\mathfrak{g}}$. It is easy to verify that this is a twisted MC element satisfying $[{\mathsf{q}},{\mathsf{q}}]=0$. Thus the commutator class vanishes in this case.
Let $G$ be a compact Lie group, $\psi$ be a gapped $G$-invariant state of a lattice system on ${\mathbb{R}}^{n}$, and $U\in{\mathcal{C}\mathcal{S}}_{n}$. The commutator class of the acyclic pointed DGLA $(C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al}),{\mathsf{Q}})$ is an obstruction for the existence of a morphism of local Lie algebras $\rho:{{\mathfrak{g}}_{al}}\rightarrow{\mathfrak{D}}^{\psi}_{al}$ which is a lift of the Lie algebra morphism ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$. Indeed, as explained in Section 5.3, if $\rho$ exists, it induces a morphism of pointed DGLAs $C^{aug}_{\bullet}({\mathfrak{U}},{\mathbb{R}}^{n};{{\mathfrak{g}}_{al}})\rightarrow C^{aug}_{\bullet}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al})$ which in turn induces a morphism in the category ${{\mathsf{pVect}}_{\mathbb{Z}}}$ which maps the commutator class of ${{\mathfrak{g}}_{al}}$ to the commutator class of ${\mathfrak{D}}^{\psi,\bf G}_{al}$. Since the former class vanishes (see Example 6.1), so must the latter.
@@ 6.2 Construction of topological invariants
The commutator class defined in the previous section is not a useful invariant of a gapped $G$-invariant state $\psi$ because it takes values in a set which itself depends on $\psi$. It is also not invariant under LGAs (defined in Section 3.2) and thus is not an invariant of a gapped phase (see Remark 3.2). In this section we define a pairing between the commutator classes of ${\mathfrak{D}}^{\psi,\bf G}_{al}$ and spherical CS cohomology classes of the sphere at infinity $S^{n-1}$, which takes values in the algebra of $G$-invariant symmetric polynomials on ${\mathfrak{g}}$. This gives a useful invariant of a gapped phase which is also an obstruction to promoting the global symmetry $G$ of the state to a local symmetry. Additionally, we will show that the invariant does not depends on the choice of the cover ${\mathfrak{U}}$ and thus is essentially unique.
Keeping in view further generalizations, we work over a sub-site ${{{\mathcal{C}\mathcal{S}}_{n}}/W}$ [^11] This is the so-called overcategory whose objects are equipped with a morphism to $W$. Due to coflasqueness, this amounts to a restriction to objects fuzzily included in $W$. which depends on an arbitray $W\in{\mathcal{C}\mathcal{S}}_{n}$. Let ${\hat{W}}\in{\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$ be the image of $W$ under the equivalence between ${\mathcal{C}\mathcal{S}}_{n}$ and ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. We continue to denote by ${\mathfrak{D}}_{al}$ the local Fréchet-Lie algebra which maps $U\in{{{\mathcal{C}\mathcal{S}}_{n}}/W}$ to ${\mathfrak{D}}_{al}$. For any cover ${\mathfrak{U}}$ of $W$ we denote by $\hat{\mathfrak{U}}$ the corresponding cover of $\hat{W}$.
Definition 6.3.
For any ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ covering $W\in{\mathcal{C}\mathcal{S}}_{n}$, any ${\mathsf{f}}\in C_{p}({\mathfrak{U}},W;{\mathfrak{D}}_{al})$, and any $\beta\in\check{C}^{p}(\hat{\mathfrak{U}},\hat{W};{\mathbb{R}})$ define an evaluation
$\langle{\mathsf{f}},\beta\rangle=\sum_{s\in I_{p}}\beta_{s}{\mathsf{f}}_{s}\in{\mathfrak{D}}_{al}(W),$
where $I_{k}:=\{i_{0}<i_{1}<\dots<i_{k}\in I^{k+1}\}$. We adopt the convention that $\beta_{s}=0$ for $\bigwedge_{j\in s}U_{j}$ bounded. The definition implicitly uses co-restriction of the coflasque cosheaf.
Lemma 6.1.
For any ${\mathfrak{U}}$ covering $W\in{\mathcal{C}\mathcal{S}}_{n}$, any cycle ${\mathsf{f}}\in C_{p}({\mathfrak{U}},W;{\mathfrak{D}}_{al})$, and any cocycle $\beta\in\check{C}^{p}(\hat{\mathfrak{U}},\hat{W};{\mathbb{R}})$ the derivation $\langle{\mathsf{f}},\beta\rangle\in{\mathfrak{D}}_{al}(W)$ is inner.
@@
The chain complex
$\ldots\rightarrow\bigoplus_{i<j}{\mathfrak{D}}_{al}(U_{i}\wedge U_{j})\rightarrow\bigoplus_{i}{\mathfrak{D}}_{al}(U_{i})\rightarrow{\mathfrak{D}}_{al}(W)\rightarrow 0$ (23)
is acyclic. Since ${\mathsf{f}}$ is a cycle, ${\mathsf{f}}=\partial{\mathsf{g}}$ for some ${\mathsf{g}}\in\bigoplus_{s\in I_{p+1}}{\mathfrak{D}}_{al}(\bigwedge_{j\in s}U_{j}).$ Thus
@@
$=$ $\sum_{s\in I_{p+1}}{\mathsf{g}}_{s}(\partial^{*}\beta)_{s}$
$=$ $\sum_{\begin{subarray}{c}s\in I_{p+1},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{bounded}\end{subarray}}{\mathsf{g}}_{s}(\partial^{*}\beta)_{s},$
@@
$(\partial^{*}\beta)_{i_{0}\dots i_{p}}=\sum_{k=0}^{p}(-1)^{k}\beta_{i_{0}\dots\widehat{i_{k}}\dots i_{p}}.$
The last expression is clearly inner because each ${\mathsf{g}}_{s}$ is almost localized near some bounded region. The last equality depends on vanishing of
$\sum_{\begin{subarray}{c}s\in I_{p+1},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{unbounded}\end{subarray}}{\mathsf{g}}_{s}(\partial^{*}\beta)_{s},$
because when $\bigwedge_{j\in s}U_{j}$ is unbounded $(\partial^{*}\beta)_{s}=(\check{\delta}\beta)_{s}=0$ where $\check{\delta}$ is the coboundary of $\check{C}^{p}(\hat{\mathfrak{U}},\hat{W};{\mathbb{R}})$. ∎
Remark 6.1.
The above lemma remains true if one replaces ${\mathfrak{D}}_{al}$ with ${\mathfrak{D}}_{al}\otimes A$ where $A$ is any locally-finite graded vector space. Then the pairing $\langle{\mathsf{f}},\beta\rangle$ takes values in the space of inner derivations ${{\mathfrak{d}}_{al}}$ (Definition 3.3) tensored with $A$.
Remark 6.2.
The relation between $\partial^{*}$ and $\check{\delta}$ is as follows. They are equal for $s\in I_{p+1}$ with $\bigwedge_{j\in s}U_{j}$ unbounded. For $s$ with $\bigwedge_{j\in s}U_{j}$ bounded, $\bigcap_{j\in s}\hat{U}_{j}=\varnothing$ and $(\partial^{*}\beta)_{s}$ is in general nonzero whereas $(\check{\delta}\beta)_{s}=0$. This discrepancy arises because the Grothendieck topology on the poset of spherical CS sets induced by the coherent topology on ${\mathcal{C}\mathcal{S}}_{n}$ does not allow the empty cover of the empty set. It is this discrepancy that makes the evaluation of DGLA cycles on spherical CS cocycles non-trivial.
If $U,V$ are closed semilinear sets with a bounded meet then any derivation in $[{\mathfrak{D}}_{al}(U),{\mathfrak{D}}_{al}(V)]$ is inner and thus has a well-defined average in any state on ${\mathscr{A}}$. More generally, if $A$ is a locally-finite supercommutative graded algebra and ${\mathfrak{F}}={\mathfrak{D}}_{al}\otimes A$ or some sub-algebra thereof, any element of $[{\mathfrak{F}}(U),{\mathfrak{F}}(V)]$ for $U\wedge V$ bounded has a well-defined average in any state of ${\mathscr{A}}$. The average takes values in $A$. We will need the following more refined result:
Lemma 6.2.
Let $U$ and $V$ be closed semilinear sets such that $U\wedge V$ is bounded. Let $A$ be a locally-finite supercommutative graded algebra and $\psi$ be a state. Then $\psi([{\mathfrak{D}}^{\psi}_{al}(U)\otimes A,{\mathfrak{D}}_{al}(V)\otimes A])=0.$

### v3

@@ 6 Topological invariants of $G$-invariant gapped states
In this section we use the twisted Maurer-Cartan equation for a pointed DGLA to construct topological invariants of gapped $G$-invariant states. For states of quantum lattice systems on ${\mathbb{R}}^{n}$, these invariants take values in $G$-invariant polynomials on the Lie algebra of $G$. We show that the Hall conductance of a $U(1)$-invariant gapped lattice system on ${\mathbb{R}}^{2}$ is a special case of this construction. We also discuss how to extend the theory to quantum lattice systems on a polyhedral subset of ${\mathbb{R}}^{n}$. In that case, topological invariants take values in $G$-invariant polynomials on the Lie algebra of $G$ valued in the cohomology of the corresponding spherical polyhedral subset of $S^{n-1}$ (“the sphere at infinity”).
@@ 6.1 The commutator class
Even if $\mathfrak{M}$} is acyclic, the DGLA $\overline{[{\mathfrak{M}},{\mathfrak{M}}]}$} is not necessarily acyclic. We would like to construct an obstruction to finding an element ${\mathsf{p}}\in{\mathfrak{M}}_{-1}$ which satisfies $\partial{\mathsf{p}}={\mathsf{B}}$ and ${\mathsf{p}},{\mathsf{p}}]=0$.
Definition 104.
@@
$\partial{\mathsf{p}}+\frac{1}{2}[{\mathsf{p}},{\mathsf{p}}]={\mathsf{B}}.$ (127)
@@
Let $({\mathfrak{M}},{\mathsf{B}})$ be a pronilpotent pointed DGLA and ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$. Then ${\mathsf{p}},{\mathsf{p}}]$ is a cycle of the DGLA $\overline{[{\mathfrak{M}},{\mathfrak{M}}]}$}.
Proposition 105.
Let $({\mathfrak{M}},{\mathsf{B}})$ be an acyclic pointed DGLA which is a limit of an inverse system of nilpotent acyclic pointed DGLAs $({\mathfrak{M}}_{N},{\mathsf{B}}_{N})$, $N\in{\mathbb{N}}$. Assume further that the structure morphisms $r_{N,N-1}:{\mathfrak{M}}_{N}\rightarrow{\mathfrak{M}}_{N-1}$ are surjective and ${\mathfrak{j}}_{N}=\ker r_{N,N-1}$ is central in ${\mathfrak{M}}_{N}$. Finally, assume ${\mathfrak{M}}_{1},{\mathfrak{M}}_{1}]=0$. Then ${\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is non-empty. Furthermore, the homology class of ${\mathsf{p}},{\mathsf{p}}]$ in $\overline{[{\mathfrak{M}},{\mathfrak{M}}]}$} for ${\mathsf{p}}\in{\rm MC}({\mathfrak{M}},{\mathsf{B}})$ is independent of the choice of $\mathsf{p}$}.
A proof of this result can be found in Appendix B. It uses some results from deformation theory. We will call the homology class of ${\mathsf{p}},{\mathsf{p}}]$ the commutator class of the acyclic pointed DGLA $({\mathfrak{M}},{\mathsf{B}})$, for lack of a better name, and denote it ${\mathbf{com}}({\mathfrak{M}},{\mathsf{B}})$. The assignment $({\mathfrak{M}},{\mathsf{B}})\mapsto(H_{\bullet}({\overline{[{\mathfrak{M}},{\mathfrak{M}}]}}),{\mathbf{com}}({\mathfrak{M}},{\mathsf{B}}))$ is a functor from the full sub-category of the category of acyclic pointed DGLAs satisfying the conditions of Proposition 105, to the category $\mathsf{pVect}$$_{\mathbb{Z}}$} of pointed graded vector spaces.
The conditions of Prop. 105 apply to any pointed DGLA which is the value of the Čech functor on a graded local Lie system ${\mathfrak{F}}^{\bf G}$ over some $(X,\leq)$, where $\mathfrak{F}$} is a local Lie system equipped with a smooth infinitesimally inner $G$-action. Indeed, along with the graded algebra $A=\prod_{k=1}^{\infty}{\rm Sym}^{k}({\mathfrak{g}}^{*}[-2])$ used in the construction of ${\mathfrak{F}}^{\bf G}$ one can consider its quotient by the ideal $J_{N}=\prod_{k=N+1}^{\infty}{\rm Sym}^{k}({\mathfrak{g}}^{*}[-2])$. Replacing $A$ with the nilpotent graded algebra $A/J_{N}$ throughout, for any cover $\mathfrak{U}$} of the terminal object $T$ one gets a sequence of nilpotent acyclic pointed DGLAs labeled by $N\in{\mathbb{N}}$. They assemble into an inverse system in an obvious manner, and its limit is the acyclic pointed DGLA $C^{aug}_{\bullet+1}({\mathfrak{U}},T;{\mathfrak{F}}^{\bf G}).$ It is easy to see that the remaining conditions of Prop. 105 are also satisfied. In particular, Prop. 105 applies to the pointed DGLAs associated to the graded local Lie systems ${\mathfrak{D}}^{\psi,\bf G}_{al}$ and ${\mathfrak{g}}^{\bf G}_{al}$} and any CS cover of ${\mathbb{R}}^{n}$.
Proposition 106.
@@
Let ${\mathfrak{V}}=\{V_{j}\}_{j\in J}$ be a refinement of $\mathfrak{U}$}. By definition, there exists a map $\phi:J\rightarrow I$ with $V_{j}\leq U_{\phi(j)}.$ According to Prop. 62, there is a map from ${\mathfrak{M}}({\mathfrak{V}})$ to ${\mathfrak{M}}({\mathfrak{U}})$. As $\mathsf{B}$} is unaffected by $\phi_{*}$, we deduce from ${\mathsf{p}}^{\mathfrak{V}}\in{\rm MC}({\mathfrak{M}}({\mathfrak{V}}),{\mathsf{B}})$ that $\phi_{*}{\mathsf{p}}^{\mathfrak{V}}\in{\rm MC}({\mathfrak{M}}({\mathfrak{U}}),{\mathsf{B}})$. ∎
Proposition 107.
Let $G$ be a compact Lie group, $\mathfrak{g}$} be its Lie algebra, and ${\mathfrak{g}}^{\bf G}_{al}$} be the pointed graded local Lie system over ${\mathcal{C}\mathcal{S}}_{n}$ of Example (102). For any cover ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ the commutator class vanishes.
Proof.
The distinguished central cycle in $C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n},{{\mathfrak{g}}^{\bf G}_{al}})$ is the constant function on $\Lambda$ with value ${\rm id}_{\mathfrak{g}}$ . Here ${\rm id}_{\mathfrak{g}}$ is regarded as a $G$-invariant element of ${{\mathfrak{g}}_{al}}({\mathbb{R}}^{n})\otimes{\mathfrak{g}}^{*}[-2]$. For any cover ${\mathfrak{U}}=\{U_{i}\}_{i\in I}$ one can construct a twisted MC element $\mathsf{q}$} as follows. Pick $r>0$ large enough so that the interiors of $U_{i}^{r}$, $i\in I$, cover ${\mathbb{R}}^{n}$ in the usual sense and pick a partition of unity $\chi_{i}$, $i\in I$, subordinate to this open cover. For any $x\in\Lambda$ and any $i\in I$ let ${\mathsf{q}}_{i}(x)=\chi_{i}(x){\rm id}_{\mathfrak{g}}$. It is easy to verify that this is a twisted MC element satisfying ${\mathsf{q}},{\mathsf{q}}]=0$. Thus, the commutator class vanishes. ∎
Corollary 108.
Let $G$ be a compact Lie group, $\psi$ be a gapped $G$-invariant state of a lattice system on ${\mathbb{R}}^{n}$, and $\mathfrak{U}$} be a ${\mathcal{C}\mathcal{S}}_{n}$-cover of ${\mathbb{R}}^{n}$. The commutator class of the acyclic pointed DGLA $(C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al}),{\mathsf{Q}})$ is an obstruction for the existence of a morphism of local Lie systems $\rho:{{\mathfrak{g}}_{al}}\rightarrow{\mathfrak{D}}^{\psi}_{al}$ which is a lift of the Lie algebra morphism ${\mathsf{Q}}:{\mathfrak{g}}\rightarrow{\mathfrak{D}}^{\psi}_{al}({\mathbb{R}}^{n})$.
Proof.
If $\rho$ exists, it induces a morphism of pointed DGLAs $C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{{\mathfrak{g}}^{\bf G}_{al}})\rightarrow C^{aug}_{\bullet+1}({\mathfrak{U}},{\mathbb{R}}^{n};{\mathfrak{D}}^{\psi,\bf G}_{al})$ which in turn induces a morphism in the category $\mathsf{pVect}$$_{\mathbb{Z}}$} which maps the commutator class of ${\mathfrak{g}}_{al}$} to the commutator class of ${\mathfrak{D}}^{\psi,\bf G}_{al}$. Since the former class vanishes, so must the latter. ∎
@@ 6.2 Construction of topological invariants
The commutator class defined in the previous section is not a useful invariant of a gapped $G$-invariant state $\psi$ because it takes values in a set which itself depends on $\psi$ and the choice of the cover. It is also not invariant under LGAs (defined in Section 2.2) and thus is not an invariant of a gapped phase (see Remark 46). In this section we define a pairing between the commutator classes of ${\mathfrak{D}}^{\psi,\bf G}_{al}$ and spherical CS cohomology classes of the sphere at infinity $S^{n-1}$, which takes values in the algebra of $G$-invariant symmetric polynomials on $\mathfrak{g}$}. This gives a useful invariant of a gapped phase which is also an obstruction to promoting the global symmetry $G$ of the state to a local symmetry. Additionally, we will show that the invariant does not depend on the choice of the cover $\mathfrak{U}$} and is controlled by the large-scale topology.
Keeping in view a generalization to quantum lattice systems on general CS sets in ${\mathbb{R}}^{n}$, we work over a sub-site ${{\mathcal{C}\mathcal{S}}_{n}}/W$} [^13] This is the so-called overcategory whose objects are equipped with a morphism to $W$. Due to coflasqueness, this amounts to a restriction to objects fuzzily included in $W$. which depends on an arbitrary $W\in{\mathcal{C}\mathcal{S}}_{n}$. Let ${\widehat{W}}\in{\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$ be the image of $W$ under the equivalence between ${\mathcal{C}\mathcal{S}}_{n}$ and ${\mathcal{S}\mathcal{C}\mathcal{S}}_{n}$. For any cover $\mathfrak{U}$} of $W$ we denote by $\hat{\mathfrak{U}}$ the corresponding cover of $\widehat{W}$.
Definition 109.
[…]
$\langle{\mathsf{f}},\beta\rangle=\sum_{s\in I^{p+1}}\beta_{s}{\mathsf{f}}_{s}\in{\mathfrak{F}}(W),$ (128)
We adopt the convention that $\beta_{s}=0$ for $\bigwedge_{j\in s}U_{j}$ bounded. The definition implicitly uses co-restriction of the coflasque cosheaf $\mathfrak{F}$}.
Let $\mathfrak{F}$} be a local Lie system over ${{\mathcal{C}\mathcal{S}}_{n}}/W$}. By analogy with Def. 39, we will say that ${\mathsf{a}}\in{\mathfrak{F}}(W)$ is anchored if $\mathsf{a}$} belongs to the sub-algebra ${\mathfrak{F}}(\{0\})$ (or equivalently, to ${\mathfrak{F}}(U)$ for some bounded $U$).
Lemma 110.
[…]
@@
The augmented chain complex
$\ldots\rightarrow{\rm Alt}\left[\bigoplus_{i,j\in I}{\mathfrak{F}}(U_{i}\wedge U_{j})\right]\rightarrow\bigoplus_{i\in I}{\mathfrak{F}}(U_{i})\rightarrow{\mathfrak{F}}(W)\rightarrow 0$ (129)
is acyclic. Since $\mathsf{f}$} is a cycle, ${\mathsf{f}}=\partial{\mathsf{g}}$ for some ${\mathsf{g}}\in{\rm Alt}\left[\bigoplus_{s\in I^{p+2}}{\mathfrak{F}}(\bigwedge_{j\in s}U_{j})\right].$ Thus
@@
$=$ $\sum_{s\in I^{p+2}}{\mathsf{g}}_{s}(\partial^{*}\beta)_{s}$
$=$ $\sum_{\begin{subarray}{c}s\in I^{p+2},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{bounded}\end{subarray}}{\mathsf{g}}_{s}(\partial^{*}\beta)_{s},$ (130)
@@
$(\partial^{*}\beta)_{i_{0}\dots i_{p+1}}=\sum_{k=0}^{p+1}(-1)^{k}\beta_{i_{0}\dots\widehat{i_{k}}\dots i_{p+1}}.$ (131)
The last expression is clearly anchored because each ${\mathsf{g}}_{s}$ is almost localized near some bounded region. The last equality depends on the vanishing of
$\sum_{\begin{subarray}{c}s\in I^{p+2},\ \mathrm{with}\\ \bigwedge_{j\in s}U_{j}\ \mathrm{unbounded}\end{subarray}}{\mathsf{g}}_{s}(\partial^{*}\beta)_{s},$ (132)
[…]
Remark 111.
The relation between $\partial^{*}$ and $\check{\delta}$ is as follows. They are equal for $s\in I^{p+2}$ with $\bigwedge_{j\in s}U_{j}$ unbounded. For $s$ with $\bigwedge_{j\in s}U_{j}$ bounded, $\bigcap_{j\in s}\hat{U}_{j}=\varnothing$ and $(\partial^{*}\beta)_{s}$ is, in general, nonzero whereas $(\check{\delta}\beta)_{s}=0$. This discrepancy arises because the Grothendieck topology on the poset of spherical CS sets corresponding to Definition 74 does not allow the empty cover of the empty set. It is this discrepancy that makes the evaluation of DGLA cycles on spherical CS cocycles non-trivial.
We are going to apply this lemma to the pointed graded local Lie system ${\mathfrak{D}}^{\psi,\bf G}_{al}$ arising from a $G$-invariant gapped state of a quantum lattice system on $W\subseteq{\mathbb{R}}^{n}$. We take $\mathsf{f}$} to be the commutator class ${\mathsf{p}},{\mathsf{p}}]$ defined via the twisted Maurer-Cartan equation. In this case $\Sigma\langle{\mathsf{f}},\beta\rangle$ takes values in the $G$-invariant part of ${{\mathfrak{d}}^{\psi}_{al}}\otimes A$ where $A=\prod_{k=1}^{\infty}{\rm Sym}^{k}({\mathfrak{g}}^{*}[-2])$. (Recall that $\Sigma$ is the isomorphism of the space of anchored elements of ${\mathfrak{D}}^{\psi}_{al}$ and ${\mathfrak{d}}^{\psi}_{al}$}). The final step in the construction of the topological invariant is to evaluate the expectation value $\psi(\Sigma\langle[{\mathsf{p}},{\mathsf{p}}],\beta\rangle)$. Since $\psi$ is $G$-invariant, it takes values in the complexification of $A^{G}$. The main result of this section is
Theorem 112.
[…]
$\nu_{\psi}(\beta)={\sqrt{-1}}\psi(\Sigma\langle[{\mathsf{p}},{\mathsf{p}}],\beta\rangle).$ (133)
The function $\nu_{\psi}$ vanishes on even-degree cohomology.
Lemma 113.
Let $U$ and $V$ be closed semilinear sets such that $U\wedge V$ is bounded. Let $A$ be a locally-finite supercommutative graded algebra and $\psi$ be a state. Then $\psi(\Sigma[{\mathfrak{D}}^{\psi}_{al}(U)\otimes A,{\mathfrak{D}}_{al}(V)\otimes A])=0.$